# The Lax–Wendroff theorem for Patankar-type methods applied to hyperbolic conservation laws

### v1

@@ 4.1 Extension to Patankar-type RK
The total time variation of a numerical solution $\{U_{i}^{n}\}_{i,n}=\{U(x_{i},t_{n})\}_{i,n}$ obtained by an s-stage RK method (TTVRK) over a domain $[a,b]\times[0,T]$ for $i=1,\dots,N$ and $n=0,\dots,n_{T}$ with $x_{1}=a$, $x_{N}=b$, $t^{0}=0$ and $t^{n_{T}}=T$ is
$\text{TTVRK}(\bm{U}(x,[0,T])):=\sum_{n=0}^{n_{T}}\sum_{k=1}^{s}|U^{(k)}(x,t_{n})-U^{(k-1)}(x,t_{n})|+\sum_{n=1}^{n_{T}}|U(x,t_{n})-U^{(s)}(x,t_{n-1})|,$ (19)
@@
[…]
Proof 3.
@@
$\delta^{\{l\}}(x,t):=\frac{U^{\{l\}}(x,t+\Delta t)}{\sigma^{\{l\}}(x,t)}=1+\frac{U^{\{l\}}(x,t+\Delta t)-\sigma^{\{l\}}(x,t)}{\sigma^{\{l\}}(x,t)}.$
Now, we can follow the same reasoning as in the proof of Lemma 2 under usage of the TTVRK property and the convergence of the numerical solutions to a weak solution $u$ in $L^{1}_{\text{loc}}$, to show the convergence of $U^{\{l\}}(x,t+\Delta t)$ to $U^{\{l\}}(x,t)$. Furthermore we assumed $\sigma^{\{l\}}(x,t)\rightarrow U(x,t)$ which finally leads to $\lim_{l\to\infty}U^{\{l\}}(x,t+\Delta t)-\sigma^{\{l\}}(x,t)=0$ and thus to $\delta^{\{l\}}(x,t)\rightarrow 1$. We proceed considering the weights of the stages
$\gamma^{(k),\{l\}}(x,t)=\frac{U^{(k),\{l\}}(x,t)}{\pi^{(k),\{l\}}(x,t)}=1+\frac{U^{(k),\{l\}}(x,t)-\pi^{(k),\{l\}}(x,t)}{\pi^{(k),\{l\}}(x,t)}$
and see the desired results under the same arguments as in the first part of the proof.
@@
then the PWDs fulfill the extended Patankar weights convergence of Lemma 4.
Proof 4.
The needed convergence $U_{i}^{(k),\{l\}}\rightarrow U_{i}^{n}$ for $l\rightarrow\infty$ is secured by the TTVD property.
@@
Example 2 (MPRK22($\alpha$)).
Consider the two stage second-order family of modified Patankar–Runge–Kutta methods introduced in [21] and denoted by MPRK22($\alpha$), where $\alpha\geq\frac{1}{2}$ is a free parameter and the PWDs are given by $\pi_{i}^{(2)}=U_{i}^{n}$ and $\sigma_{i}=\left(U_{i}^{n}\right)^{1-\frac{1}{\alpha}}\left(U_{i}^{(2)}\right)^{\frac{1}{\alpha}}$.
@@
$\left\lvert F_{g}^{\delta,\{l\}}\left(U^{(r),\{l\}}(\mathbf{x}_{i},t_{n}^{(r)})\right)-f(U^{\{l\}}(x_{i},t_{n}))\right\rvert\leq$ $\left\lvert\max\{g(U^{(r),\{l\}}(\mathbf{x}_{i},t_{n}^{(r)})),0\}\delta_{i}^{\{l\}}-\max\{g(U^{(r),\{l\}}(\mathbf{x}_{i},t_{n}^{(r)})),0\}\right\rvert$
@@
Note in this case the RK stages are arguments of the numerical flux. Therefore, we need first Lemma 4 to secure the convergence of the stages to the same solution $\lim_{l\to\infty}U^{(r),\{l\}}(x_{i}+j\Delta x,t_{n}^{(r)})-U^{\{l\}}(x_{i}+j\Delta x,t_{n})=0$ and proceed with the proof of the original Lax–Wendroff theorem to obtain the desired result.
@@
almost everywhere. This property represents the main ingredient in the proof outlined in Remark 8. Indeed, to extend the Lax–Wendroff theorem also to general NSARK schemes with general solution-dependent coefficients of the form (8), it is sufficient to add the condition (20) to the classical assumptions of the Lax–Wendroff theorem.
Remark 10 (On the TTV boundness assumption).
The difference with respect the classical Lax–Wendroff theorem consists on the assumption that the numerical solutions are TTV bounded uniformly in space and in the limit process. The authors could not easily remove this assumption to prove the Lax–Wendroff theorem. Different strategies have been pursued but all of them ended up requiring some extra hypotheses. In particular, the technique proposed by Birken and Linders [6] shows that Lax–Wendroff theorem using iterative methods for implicit schemes can be applied also in the context of Patankar-type methods. Nevertheless, in their work the case where both the limit of iterations of the linear solver and the mesh refinement level go to infinity is not treated. In particular, it is hard to prove the $\lim_{l\to\infty}\lim_{z\to\infty}$ case, where $l$ is the mesh refinement level and $z$ is the linear solver iteration. Indeed, this leads to a dependency of the composed numerical flux in one step of the RK on the whole domain. This destroys the argument used in the classical Lax–Wendroff theorem to prove the flux convergence almost everywhere.
@@ 5 Numerical results
In this section, we will run some simple Riemann problem tests to validate the convergence towards exact solutions. What we are looking for is to test the regimes where the modified Patankar weights are large and change significantly the scheme with respect to explicit ones. Clearly, this happens close to discontinuity and close to 0 values of the solution, while everywhere else the modified Patankar weights quickly converges to 1. Hence, we choose to test Riemann problems with very low value of one of the two sides.
@@ 5.1.1 Burgers’ equation
$u(x,0)=\begin{cases}u_{1},&\text{ if }-0.5<x<0.5\\ u_{2},&\text{ else},\end{cases}$ (22)
@@
To check the convergence of the scheme, we look at a very strong Riemann problem where $u_{1}=10^{4}$ and $u_{2}=10^{-30}$, so that we expect the modified Patankar weights to be particularly large. We run some simulations with different number of cells $N$ at CFL = 1 and CFL = 2.1. For each of these simulations we look for the interface with largest derivative and we set this point to be the “numerical shock location” $s_{N}$. We measure then $|s_{N}-s|$, where $s$ is the exact shock location. For both CFLs (so, also in a non TVD regime), we obtain that the numerical shock locations converge linearly to the exact shock location.
@@ 5.1.2 Buckley–Leverett equation
In Figure 2, we show the simulations of explicit Euler and MPE at various $N$ for $\text{CFL}=0.99$ (left), $\text{CFL}=1.2$ (center) and $\text{CFL}=1.99$ (right). Below each figure, we also plot the error of the numerical shock location. MPE clearly has some qualitative improvements with respect to explicit Euler, in particular for $\text{CFL}=1.2$ where we are outside the stability region for the explicit method, but we can still obtain reliable simulations for MPE. In the explicit method, we observe not only oscillations close to the shock, but also a wrong weak solution at the rarefaction wave (not the entropy solution), while the MPE gives correct results at both waves. Numerically, we have seen that for MPE seems TVD up to $\text{CFL}=1$, while explicit Euler is not TVD for CFL larger than 1. In any case, in the regime where explicit Euler converges to the exact solution, we also have that the MPE converges to the exact solution.
@@ 5.2.1 High order in time
In figure (4), we check the TTV boundness for different time discretization methods at different CFL number (0.99, 1.99, 2.99, 10) for a double RP (22) with $u_{1}=2$ and $u_{2}=10^{-13}$. The TTV boundness is the hypothesis that we added to the Lax–Wendroff theorem in order to be proven for Patankar-type methods. Here, we check whether this hypothesis is reasonable for such methods. In particular, we observe that the Modified Patankar methods seem to have bounded TTV also for very large CFLs, even if the bound is larger than the exact solution one (which is bounded by 2). This does not prove that the method is uniformly bounded for large CFLs, but it gives good hopes, while for low CFLs it seems that the TTV is bounded by the exact TTV. On the other side, the explicit Euler method is TTV only for $\text{CFL}\leq 1$, while for larger CFLs has an exploding behavior.
@@ 5.2.2 High order in space and time with WENO-FV
Finally, in this section we propose a space-time high order discretization which involves a finite volume (FV) WENO spatial reconstruction [40] in each cell, with a positivity limiter [36] to guarantee the positivity in each reconstruction point. Typically these limiters are set to work with the explicit Euler method and SSPRK methods, where the CLF has to be adjusted according to the weights of the underlying quadrature, e.g. for order 5 one needs a CFL of $\frac{1}{12}$. In the context of Patankar-type schemes, it has been shown that this restriction of the CFL is not required [10].
@@
We start with the inviscid Burgers’ equations $\partial_{t}u(x,t)+\partial_{x}(u^{2}(x,t)/2)=0$ for $u:[0,1]\times\mathbb{R}^{+}\to\mathbb{R}$ with IC and exact solution
@@
The solution of this problem is clearly positive, hence it is reasonable to apply the MP procedure on the variable $u$. In Fig. 5 (top), we show the simulations for various methods at time $T=1$ obtained with $\text{CFL}=0.5$. In all schemes, we see a convergence of the shock towards the exact solution, in accordance with Remark 8 and, hence, with the shock traveling at the right speed. This is a symptom of a conservative method and consistent numerical fluxes.
@@
In Fig. 5 (bottom), we depict the simulation at time $T=0.7$ obtained with $\text{CFL}=0.5$ for different MP solvers next to the exact solutions as well as the numerical shock location error. Here, we also observe that the solutions converge to the exact one as the mesh is refined.
@@ 6 Conclusions and perspectives
It would be interesting to extend this result, for example, proving that the extra hypothesis is not needed. In this context, ideas similar to [6] could be explored, even if there is no guarantee that the two limits of linear solver iterations going to infinity and mesh refinements can be exchanged. Another aspect of interest is a formulation of entropy preserving Patankar-type methods which is ongoing work.
@@ Appendix A Total Variation Diminishing property
In this section, we will prove an interesting results for a very specific combination of space-time discretization on Burgers’ equation. We use MPE as time discretization, an upwind numerical flux (being $u$ positive, this is quite straightforward) with the definition of PDS of (9). We can prove that this scheme is total variation diminishing (TVD) for an implicit $\text{CFL}<2$.

### v2

@@ 4.1 Extension to Patankar-type RK
The total time variation of a numerical solution $\{U_{i}^{n}\}_{i,n}=\{U(x_{i},t_{n})\}_{i,n}$ obtained by an s-stage RK method (TTVRK) over a domain $a,b]\times[0,T]$ for $i=1,\dots,N$ and $n=0,\dots,n_{T}$ with $x_{1}=a$, $x_{N}=b$, $t_{0}=0$ and $t_{n_{T}}=T$ is
$\text{TTVRK}(\bm{U}(x,[0,T]))\mathrel{\mathop{\ordinarycolon}}=\sum_{n=0}^{n_{T}}\sum_{k=1}^{s}|U^{(k)}(x,t_{n})-U^{(k-1)}(x,t_{n})|+\sum_{n=1}^{n_{T}}|U(x,t_{n})-U^{(s)}(x,t_{n-1})|,$ (19)
@@
Let $U^{\{l\}}(x_{i},t_{n})$ be the approximation of problem (1) at the space $x^{\{l\}}_{i}$ and time $t_{n}^{\{l\}}$ for a mesh discretization $\Delta x_{\{l\}}\sim\Delta t_{\{l\}}\to 0$ for $l\to\infty$. For every $l$, consider $U^{\{l\}}(x,t)$ a piece-wise constant reconstruction of the numerical solution $U^{\{l\}}(x_{i},t_{n})$. Let $\text{TTVRK}(U^{\{l\}}(x,[0,T]))$ be uniformly bounded for all $x\in[a,b]$ and for all $l\in\mathbb{N}$. Suppose that $U^{\{l\}}\to u$ in $L^{1}_{\text{loc}}([a,b]\times[0,T])$. Defining $\delta^{\{l\}}(x,t)$ as the piece-wise constant reconstruction of $\delta^{\{l\}}(x_{i},t_{n})=U^{\{l\}}(x_{i},t_{n+1})/\sigma^{\{l\}}(x_{i},t_{n})$ and $\gamma^{(k),\{l\}}(x,t)$ the one of $\gamma^{(k),\{l\}}(x_{i},t_{n}^{k})=U^{(k),\{l\}}(x_{i},t_{n}^{k})/\pi^{(k),\{l\}}(x_{i},t_{n}^{k})$ for all $i$ and $n$. Then, $\delta^{\{l\}}(x,t)$ and $\gamma^{(k),\{l\}}(x,t)\to 1$ almost everywhere in $a,b]\times[0,T]$ if $\sigma^{\{l\}}(x,t)$, respectively, $\pi^{(k),\{l\}}(x,t)\rightarrow u(x,t)$ as $l\to\infty$.
Proof.
@@
$\delta^{\{l\}}(x,t)\mathrel{\mathop{\ordinarycolon}}=\frac{U^{\{l\}}(x,t+\Delta t_{\{l\}})}{\sigma^{\{l\}}(x,t)}=1+\frac{U^{\{l\}}(x,t+\Delta t_{\{l\}})-\sigma^{\{l\}}(x,t)}{\sigma^{\{l\}}(x,t)}.$
Now, we can follow the same reasoning as in the proof of Lemma 2 under usage of the uniformly bounded TTVRK and the convergence of the numerical solutions to a weak solution $u$ in $L^{1}_{\text{loc}}$, to show $\lim_{l\to\infty}U^{\{l\}}(x,t+\Delta t_{\{l\}})=u(x,t)$ a.e.. Furthermore we assumed that $\sigma^{\{l\}}(x,t)\rightarrow u(x,t)$ which finally leads to $\lim_{l\to\infty}U^{\{l\}}(x,t+\Delta t_{\{l\}})-\sigma^{\{l\}}(x,t)=0$ and thus to $\delta^{\{l\}}(x,t)\rightarrow 1$. We proceed considering the weights of the stages
$\gamma^{(k),\{l\}}(x,t)=\frac{U^{(k),\{l\}}(x,t)}{\pi^{(k),\{l\}}(x,t)}=1+\frac{U^{(k),\{l\}}(x,t)-\pi^{(k),\{l\}}(x,t)}{\pi^{(k),\{l\}}(x,t)}.$
Once again, we use the fact that TTVRK is uniformly bounded, together with the convergence to a weak solution $u$ for the limit $\lim_{l\to\infty}U^{(k),\{l\}}(x,t)=u(x,t)$. Additionally, we assumed $\pi^{(k),\{l\}}(x,t)\rightarrow u(x,t)$ as $l\to\infty$, so this leads to $\lim_{l\to\infty}U^{(k),\{l\}}(x,t)-\pi^{(k),\{l\}}(x,t)=0$ and thus to the desired result. ∎
@@
Then, the PWDs fulfill the extended Patankar weights convergence of Lemma 4.
Proof.
Recall that $U^{(k),\{l\}}\rightarrow u$ for $l\rightarrow\infty$ by means of the uniform boundedness of the TTVRK together with $U^{\{l\}}\to u$ in $L^{1}_{\text{loc}}([a,b]\times[0,T])$, as mentioned in the proof of Lemma 4. Hence, the arguments of the PDWs tend to $u$, and thus, the continuity and consistency of the PWDs yield $\sigma^{\{l\}}(x,t),\,\pi^{(k),\{l\}}(x,t)\rightarrow u(x,t)$ almost everywhere for $l\rightarrow\infty$ and $(x,t)\in[a,b]\times[0,T]$. ∎
@@
Example 2 (Verification of PWD Hypothesis for MPRK22($\alpha$)).
Consider the two stage second-order family of MPRK methods introduced in [23] and denoted by MPRK22($\alpha$), where $\alpha\geq\frac{1}{2}$ is a free parameter and the PWDs are given by $\pi_{i}^{(2)}=U_{i}^{n}$ and $\sigma_{i}=\left(U_{i}^{n}\right)^{1-\frac{1}{\alpha}}\left(U_{i}^{(2)}\right)^{\frac{1}{\alpha}}$.
@@
[…]
@@
Note in this case the RK stages are arguments of the numerical flux. Therefore, we first need the TTVRK boundedness property to apply Lemma 4 and ensure that the stages converge to the same solution $\lim_{l\to\infty}U^{(r),\{l\}}(x_{i}+j\Delta x,t_{n}^{(r)})-U^{\{l\}}(x_{i}+j\Delta x,t_{n})=0$ and proceed with the proof of the original Lax–Wendroff theorem to obtain the desired result.
@@
almost everywhere. This property represents the main ingredient in the proof outlined in Remark 9. Indeed, to extend the Lax–Wendroff theorem also to general NSARK schemes with general solution-dependent coefficients of the form (8), it is sufficient to add the condition (20) to the classical assumptions of the Lax–Wendroff theorem. In the case of MPRK schemes, we used the TTVRK boundedness together with special properties of the PWDs (see Theorem 5) to deduce (20) in Lemma 4.
Remark 11 (On the TTV boundedness assumption).
The difference with respect to the classical Lax–Wendroff theorem consists of the assumption that the numerical solutions are TTV bounded uniformly in space and in the limit process. The authors could not easily remove this assumption to prove the Lax–Wendroff theorem. Different strategies have been pursued but all of them ended up requiring some extra hypotheses. In particular, the technique proposed by Birken and Linders [8] shows that, using iterative methods for implicit schemes, the Lax–Wendroff theorem can be applied also in the context of Patankar-type methods. Nevertheless, in their work the case where both the limit of iterations of the linear solver and the mesh refinement level go to infinity is not treated. In particular, it is hard to prove the $\lim_{l\to\infty}\lim_{z\to\infty}$ case, where $l$ is the mesh refinement level and $z$ is the linear solver iteration. Indeed, this leads to a dependency of the composed numerical flux in one step of the RK on the whole domain. This destroys the argument used in the classical Lax–Wendroff theorem to prove the flux convergence almost everywhere.
@@ 5 Numerical results
In this section, we will run some simple Riemann problem tests to validate the convergence towards exact solutions. What we are looking for is to test the regimes where the MP weights are large and change significantly the scheme with respect to explicit ones. Clearly, this happens close to discontinuity and close to 0 values of the solution, while everywhere else the MP weights quickly converges to 1. Hence, we choose to test Riemann problems with very low value of one of the two sides.
@@ 5.1.1 Burgers’ equation
$u(x,0)=\begin{cases}u_{1},&\text{ if }-0.5<x<0.5,\\ u_{2},&\text{ else},\end{cases}$ (22)
@@
To check the convergence of the scheme, we look at a very strong Riemann problem where $u_{1}=10^{4}$ and $u_{2}=10^{-30}$, so that we expect the MP weights to be particularly large. We run some simulations with different number of cells $N$ at CFL = 1 and CFL = 2.1. For each of these simulations we look for the interface with largest derivative and we set this point to be the “numerical shock location” $s_{N}$. We measure then $|s_{N}-s|$, where $s$ is the exact shock location. For both CFLs (so, also in a non TVD regime), we obtain that the numerical shock locations converge linearly to the exact shock location.
@@ 5.1.2 Buckley–Leverett equation
In Figure 2, we show the simulations of explicit Euler and MPE at various $N$ for $\text{CFL}=0.99$ (left), $\text{CFL}=1.2$ (center) and $\text{CFL}=1.99$ (right). Below each figure, we also plot the error of the numerical shock location. MPE clearly has some qualitative improvements with respect to explicit Euler, in particular for $\text{CFL}=1.2$ where we are outside the stability region for the explicit method, but we can still obtain reliable simulations for MPE. In the explicit method, we observe not only oscillations close to the shock, but also a wrong weak solution at the rarefaction wave (not the entropy solution), while the MPE gives correct results at both waves. Numerically, we have seen that MPE seems to be TVD up to $\text{CFL}=1$, while explicit Euler is not TVD for CFL larger than 1. In any case, in the regime where explicit Euler converges to the exact solution, we also have that the MPE converges to the exact solution.
@@ 5.2.1 High order in time
In figure (4), we check the TTV boundedness for different time discretization methods at different CFL number (0.99, 1.99, 2.99, 10) for a double RP (22) with $u_{1}=2$ and $u_{2}=10^{-13}$. The TTV boundedness is the hypothesis that we added to the Lax–Wendroff theorem in order to be proven for Patankar-type methods. Here, we check whether this hypothesis is reasonable for such methods. In particular, we observe that the MP methods seem to have bounded TTV also for very large CFLs, even if the bound is larger than the exact solution one (which is bounded by 2). This does not prove that the method is uniformly bounded for large CFLs, but it gives good hopes, while for low CFLs it seems that the TTV is bounded by the exact TTV. On the other side, the explicit Euler method is TTV only for $\text{CFL}\leq 1$, while for larger CFLs has an exploding behavior.
@@ 5.2.2 High order in space and time with WENO-FV
Finally, in this section we propose a space-time high order discretization which involves a FV WENO spatial reconstruction [42] in each cell, with a positivity limiter [38] to guarantee the positivity in each reconstruction point. Typically these limiters are set to work with the explicit Euler method and SSPRK methods, where the CFL has to be adjusted according to the weights of the underlying quadrature, e.g. for order 5 one needs a CFL of $\frac{1}{12}$. In the context of Patankar-type schemes, it has been shown that this restriction of the CFL is not required [12].
@@
We start with the inviscid Burgers’ equations $\partial_{t}u(x,t)+\partial_{x}(u^{2}(x,t)/2)=0$ for $u\mathrel{\mathop{\ordinarycolon}}[0,1]\times\mathbb{R}^{+}\to\mathbb{R}$ with IC and exact solution
@@
The solution of this problem is clearly positive, hence it is reasonable to apply the MP procedure on the variable $u$. In Fig. 5 (top), we show the simulations for various methods at time $T=1$ obtained with $\text{CFL}=0.99$. In all schemes, we see a convergence of the shock towards the exact solution, in accordance with Remark 9 and, hence, with the shock traveling at the right speed. This is a symptom of a conservative method and consistent numerical fluxes.
@@
In Fig. 5 (bottom), we depict the simulation at time $T=0.7$ obtained with $\text{CFL}=0.99$ for different MP solvers next to the exact solutions as well as the numerical shock location error. Here, we also observe that the solutions converge to the exact one as the mesh is refined.
@@ 6 Conclusions and perspectives
It would be interesting to extend this result, for example, proving that the extra hypothesis is not needed. In this context, ideas similar to [8] could be explored, even if there is no guarantee that the two limits of linear solver iterations going to infinity and mesh refinements can be exchanged. Some of the authors are also working on defining an entropy-preserving version of Patankar-type methods. This will be of particular interest for the development of Patankar-type methods that guarantee the positivity of quantities that are not directly evolved by the conservation law but instead are derived by an additional equation of state. For instance, the pressure in the context of the compressible Euler equations would be positive if the entropy is dissipated along the simulation and the density is positive, see for example [7]. This still requires some additional work and it is not a trivial modification of Patankar-type schemes for PDS.
@@ Appendix A Total Variation Diminishing property
In this section, we will prove an interesting result for a very specific combination of space-time discretization on Burgers’ equation. We use MPE as time discretization, an upwind numerical flux (being $u$ positive, this is quite straightforward) with the definition of PDS of (9). We can prove that this scheme is total variation diminishing (TVD) for an implicit $\text{CFL}<2$.
@@ A.1 Numerical results
During the preparation of this work the authors used ChatGPT (OpenAI) in order to refer to the compactness of functions of bounded variation (Lemma 2). After using this tool/service, the author(s) reviewed and edited the content as needed and take full responsibility for the content of the published article.